Suppose a security pays $100 in 6 months, and is currently trading at $96.99. What is the annual percentage rate on this security?

Round your answer to 4 decimal places. For example if your answer is 3.205%, then please write down 0.0321.
Answer #1

Expected Future Payout = $ 100, Tenure = 6 months and Current Price = $ 96.99

Let the 6-month rate be R

Therefore, 96.99 = 100 / (1+R)

R = (100 / 96.99) - 1 = 0.03103

Annual Percentage Rate (APR) = 2 x R = 2 x 0.03103 = 0.062068 ~ 0.06207